### v1

@@ Abstract
We developed a tunnel diode oscillator and characterized its performance, highlighting its potential applications in the quantum state readout of electrons in semiconductors and electrons on liquid helium. This cryogenic microwave source demonstrates significant scalability potential for large-scale qubit readout systems due to its compact design and low power consumption of only 1 $\mu$W, making it suitable for integration on the 10 mK stage of a dilution refrigerator. The tunnel diode oscillator exhibits superior amplitude stability compared to commercial microwave sources. The output frequency is centered around 140 MHz, commonly used for qubit readout of electrons in semiconductors, with a frequency tunability of 20 MHz achieved using a varactor diode. Furthermore, the phase noise was significantly improved by replacing the commercially available voltage source with a lead-acid battery, achieving a measured phase noise of -115 dBc/Hz at a 1 MHz offset.
@@ 1 Introduction
One of the key features required to realize a fault-tolerant scalable quantum computer is the integration of energy-efficient and reliable qubit control and readout electronics. Recently, qubit control electronics have been successfully integrated using cryogenic Complementary Metal-Oxide-Semiconductor (CMOS) technology [1, 2, 3, 4, 5] , and superconducting Josephson-junction [6]. Here, we focus on developing the readout electronics using tunnel-diode oscillator [7, 8] (TDO) circuits. TDO uses a tunnel diode as a negative resistance element [9], generating a microwave signal when connected to an LC circuit. Comparable to cryogenic CMOS devices (typical power consumption: 10 mW [2, 3, 4, 10, 11, 5]) and superconducting Josephson-junction circuits (100 $\mu$[6]), the TDO presented here exhibits significantly lower power dissipation, consuming only 1 $\mu$W.
@@
The motivation of this research is to develop a technique that enables scalable qubit read-out schemes for quantum computers operating at cryogenic temperatures. The conventional method, illustrated in Fig. 1(a), is to generate a microwave signal with a room-temperature (RT) microwave source and send it to a resonator dispersively coupled to the qubit at 10 mK. The amplitude and phase of the reflected microwave signal depend on the qubit state, allowing us to determine the qubit state through measurement. This method is commonly used for qubits made from electrons in semiconductors [12, 13, 14, 15, 16] and a similar method has been proposed for electrons on helium [17].
@@
Figure 1: (a, b) Schematic diagrams illustrating dispersive qubit readout using a resonator: the conventional method in (a) and the tunnel-diode oscillator (TDO) method in (b). The bandwidth of the twisted pair of DC lines is limited to 1 MHz but is sufficient for transmitting the final calculation results to room temperature (RT) and powering the TDO from RT. (c) Schematic of the setup used to characterize the TDO. The characterization includes data measured both with and without the amplifier at 4 K.
@@
As the input signal is generated at RT, attenuators are placed at 4K to reduce thermal noise from RT reaching the qubit. To prevent the reflected output signal from being overwhelmed by thermal noise at RT, it must be amplified at cryogenic temperatures. Cryogenic amplifiers typically have a lower noise figure, resulting in a higher signal-to-noise ratio. The resonant frequency of the resonator is typically in the range of 100 MHz to 2 GHz [18, 19, 15], and such high-frequency signals are transmitted through coaxial cables. A major scalability issue arises as the number of qubits increases because each qubit requires its own coaxial cable, which are relatively thick (diameter $\sim$1 cm) with attenuators and amplifiers also taking up the limited space inside a cryogenic refrigerator (diameter $\sim$10-100 cm). Multiplexing can reduce the number of coaxial cables to some extent, but is limited by the resonator bandwidth [20].
To address this challenge, we propose placing the microwave source closer to the qubits. By integrating the microwave source and qubits on the same board, we can replace bulky coaxial cables with compact on-board transmission lines. This integration significantly reduces the circuit’s size, making it more suitable for scaling up the number of qubits. Since the qubits must be operated at cryogenic temperatures (typically 10 mK), the microwave source must also function at the same temperatures. In this regard, the TDO we developed operates at the same temperature as the qubit, which gives it an advantage over cryogenic CMOS devices which should be placed at 4 K due to its high power consumption [1, 2, 3, 4, 5]. CMOS technology offers the advantage of high output power; however, such high power is not necessary for qubit readout applications.
The TDO we developed is compact, with a size of approximately $2\,\text{cm}\times 2\,\text{cm}$ (Fig. 2(a)), and could be further miniaturized. It is powered by a DC voltage supplied through DC lines. For these scenarios, an analog-to-digital converter (ADC) at cryogenic temperatures [21] (Fig. 1(b)) can be placed to remove the need for co-axial cables on the resonator output lines as well. This approach would enable faster quantum error correction [22] feedback without the need to transmit signals to RT. Only the final computation results would need to be transmitted to RT. Placing the ADC at cryogenic temperatures also eliminates the need for amplifiers, as thermal noise at RT would no longer be a concern.
@@
In this work, we first focus on the characterization of the TDO without utilizing an ADC or a resonator coupled to qubits (Fig. 1(c)). This initial step allows us to evaluate the performance and stability of the TDO as a standalone device under cryogenic conditions, providing a foundation for future integration with qubits.
@@ 2 Tunnel Diode Oscillator
We use the commercial BD-6 Ge-backward tunnel diode (General Electric) because it has the lowest power consumption among the commercially available options. TDOs with backward tunnel diodes demonstrate frequency stability of 0.001 ppm [8, 23]. Utilizing this precision, TDOs with backward tunnel diodes have been employed in the past to investigate various physical properties, such as paramagnetic susceptibility in salts [24], penetration depth in superconductors [25, 26], and the melting point of 3He [27]. These results were obtained with long integration times using a frequency counter, whereas qubit readout demands precise measurements within shorter durations. This emphasises the need to understand the frequency and amplitude stability achievable over shorter integration times. Furthermore, the oscillation frequencies, typically around 10 MHz [24, 28, 25, 26, 29, 30], are relatively low for qubit readout applications. In the early stages of testing qubit readout, it is also desirable to have the capability to externally tune the frequency. To address this issue, we developed a TDO operating at around 140 MHz utilizing a microfabricated superconducting spiral inductor, and by incorporating a varactor diode as a variable capacitor, we achieved a frequency tunability of 20 MHz without altering the amplitude.
@@
Figure 2: (a) A photograph of the developed circuit. The pink box encloses the TDO, and the green box the resonator. The orange circle highlights a microfabricated 15-turn spiral inductor on a sapphire substrate, positioned on top of a 1.5-turn pickup coil, enabling a 10% signal to be extracted from the TDO in terms of voltage. The pickup coil and the TDO are connected via a coupler (ZEDC-15). In this study, the resonator is detuned from its resonance, behaving as an open-ended element in the circuit. An enlarged photograph of the microfabricated 15-turn spiral inductor on a sapphire substrate is also shown. The diameter of its boding pad is 176 $\mu$m. (b) Circuit diagram of the device shown in (a). See the main text for details of the circuit components. The signal output from the TDO is measured at RT using a digital oscilloscope or a spectrum analyzer via AC port I or AC port II. For comparison, a signal from a commercial microwave source is sent to the cryogenic device through AC port II, and the reflected signal is measured at RT using a digital oscilloscope or a spectrum analyzer via AC port I. (c) The I-V curve of the tunnel diode BD-6. Hysteresis is observed only during oscillation, depending on the sweep direction of the DC bias $V_{\mathrm{TD}}$ applied to the tunnel diode (indicated by arrows). When the bias is increased in the forward direction, oscillation occurs in the range of 0.04 V to 0.245 V, whereas in the reverse direction, the oscillation range is from 0.015 V to 0.04 V.
We measured 62 different BD-6 diodes and observed a wide variation in parasitic capacitance, ranging from $C_{\mathrm{TD}}=$1.3 pF to 29.4 pF, while the peak and valley points of their I-V characteristics remained relatively consistent. For high frequency oscillation, selecting a BD-6 with low parasitic capacitance is preferable. The tunnel diode used in this study has a parasitic capacitance of approximately $C_{\mathrm{TD}}=$7 pF for the voltage applied to the tunnel diode (TD) $V_{\mathrm{TD}}\approx 0.1$ V. Here, we use the microfabricated 15-turn Nb spiral coil as an inductor, which was separately measured to be 95 nH at 4 K. When a solenoid coil made of normal metal with a similar inductance value was used as the inductor, oscillation could not be observed. We deduce this was because the oscillation frequency approached the self-resonant frequency of the solenoid coil inductor. This is why we employed the micro-scale spiral inductor, as it reduces the parasitic capacitance and increases the self-resonant frequency to overcome this limitation. Furthermore, by employing a superconducting material, the Q-factor of the inductor was enhanced, enabling oscillation from the TDO at around 140 MHz. Relatively high frequencies, around 100 MHz, have been achieved using a different approach based on a toroidal LC resonator [31]. However, due to its bulky nature, this approach is unsuitable for scalable qubit applications.
@@
Fig. 2(a,b) shows the details of the developed circuit. The part enclosed by the pink line represents the TDO, which consists of the Nb spiral inductor and a variable capacitance formed by a varactor diode (MA46H201, MACOM). The varactor diode allows tuning of the LC resonator’s capacitance, thereby varying the oscillation frequency of the TDO. The varactor diode is biased via DC port 2 using a DC voltage source (Yokogawa GS200). The oscillation signal is coupled to the pickup coil with a coupling ratio of 15:1.5. In previous work [24, 28, 25, 26, 29, 30], the signal was extracted capacitively from the same line used to provide the DC bias to the tunnel diode. Instead, in this work, the signal is extracted inductively. This approach is taken to dedicate the DC bias line solely to providing the DC bias, thereby avoiding interference with filters and other components further along the line. As a result, a more stable DC bias can be supplied to the tunnel diode. The tunnel diode is biased via DC port 1 using either the Yokogawa GS200 or a lead-acid battery. After the pickup coil, the signal is sent to AC port II and measured at RT. It is also routed through a directional coupler (ZEDC-15, Mini-Circuits). The coupling from the coupling port to the input/output port of the ZEDC-15 is 20 dB, as measured separately at 4 K. The signal from the coupler is amplified at 4 K with a cryogenic amplifier (CMT-BA1, Cosmic Microwave Technology) before being sent to RT.
@@
The part enclosed by the green line in Fig. 2(a, c) represents the resonator, which can be coupled to a qubit. In this work, the resonator’s resonant frequency is set far outside the TDO’s operating range, allowing it to act as a high-impedance element and remain unused. The signal output from the coupler is sent to a cryogenic low-noise amplifier (CMT-BA1, Cosmic Microwave Technology) at 4 K and is measured at RT using a spectrum analyzer or a digital oscilloscope.
@@
As shown in Fig. 2(c), the operating point of the tunnel diode BD-6, where it exhibits negative resistance and oscillates, is approximately $\,\text{V}$ 0.1 V and $I\approx 10\,\mu\text{A}$, resulting in a power consumption of about $1\,\mu\text{W}$.
@@ 3 Frequency and power tunability
In this section, we demonstrate that the output power and frequency can be controlled by varying the voltage applied to the tunnel diode and that the frequency can be independently adjusted by varying the voltage applied to the varactor diode. The frequency variation due to the voltage applied to the tunnel diode (TD) $V_{\mathrm{TD}}$ is shown in Fig. 3(a). The capacitance of the TD, $C_{\mathrm{TD}}$, is voltage-dependent because the distance of the depletion layer in the p-n junction changes as a function of $V_{\mathrm{TD}}$. It is given by $C_{\mathrm{TD}}=\frac{C_{0}}{\sqrt{V_{\mathrm{d}}-V_{\mathrm{TD}}}}$ [7], where $V_{d}=0.5$ V is the diffusion potential (estimated from a measurement), and the equation $f=\frac{1}{2\pi\sqrt{L(C_{p}+C_{\mathrm{TD}})}}$ has been fitted, as shown by the dashed lines, where $C_{p}$ represents the sum of the parasitic capacitance of the circuit board and the capacitance of the varactor diode. The fitting results give $C_{p}=4.85$ pF and $C_{0}=4.7$ pF at 11 mK, and $C_{p}=4.94$ pF and $C_{0}=4.75$ pF at 3.4 K. Fig. 3(b) demonstrates that the output power can be adjusted by approximately 20 dB through tuning $V_{\mathrm{TD}}$ because the operating point of the tunnel diode shifts along its I-V curve. The microwave power that would be delivered to the resonator is around -90 dBm, which is comparable to the optimized power range or qubit readout of electrons used as qubits in semiconductors [32].
@@
Fig. 3(c) shows the frequency tunability as a function of $V_{\mathrm{VD}}$. Considering the additional changes due to $V_{\mathrm{TD}}$, we observe a total frequency tunability of 20 MHz. From this, we estimated that $C_{p}$ varies from 4.2 pF at $V_{\mathrm{VD}}=5$ V to 6.3 pF at $V_{\mathrm{VD}}=-1.5$ V. Fig. 3(d) shows that for $V_{\mathrm{VD}}>-1.3$ V, the output power remains unaffected by $V_{\mathrm{VD}}$, indicating that the resistive component of the varactor remains unchanged (leakage current stays significantly small within this voltage range as seen in Fig. 6). Please note that we have observed a change in output power when $V_{\mathrm{VD}}$ is varied within the same voltage range at RT when using a similar circuit with a normal coil inductor. We attribute this difference to the fact that the leakage current becomes significantly suppressed at low temperatures.
@@
Figure 3: The signal from the TDO obtained through AC port II was measured using a spectrum analyzer. (a,b) $V_{\mathrm{VD}}$ is set to 0. The sweep direction of $V_{\mathrm{TD}}$ is forward. (a) Output frequency as a function of $V_{\mathrm{TD}}$. The dashed lines represent the fits, taking into account the change in depletion layer thickness induced by varying $V_{\mathrm{TD}}$. See the main text for details on the fitting procedure. (b) Output power as a function of $V_{\mathrm{TD}}$. The difference in the background level is attributed to the use of different spectrum analyzers for measurements at 11 mK and 3.4 K. (c) Output frequency as a function of $V_{\mathrm{VD}}$. (d) Output power as a function of $V_{\mathrm{VD}}$ and remains constant for $V_{\mathrm{VD}}>-1.3$ V.

### v2

@@ Abstract
We developed a tunnel diode oscillator and characterized its performance, demonstrating its potential applications in the quantum state readout of electrons in semiconductors and electrons on liquid helium. This cryogenic microwave source demonstrates significant scalability potential for large-scale qubit readout systems due to its compact design and low power consumption of only 1 $\mu$W, making it suitable for integration on the 10 mK stage of a dilution refrigerator. The tunnel diode oscillator exhibits superior amplitude stability compared to commercial microwave sources. The output frequency is centered around 140 MHz, commonly used for qubit readout of electrons in semiconductors, with a frequency tunability of 10 MHz achieved using a varactor diode. Furthermore, the phase noise was significantly improved by replacing the commercially available voltage source with a lead-acid battery, achieving a measured phase noise of -115 dBc/Hz at a 1 MHz offset.
@@ 1 Introduction
One of the key features required to realize a fault-tolerant scalable quantum computer is the integration of energy-efficient and reliable qubit control and readout electronics. Recently, qubit control electronics have been successfully integrated using cryogenic Complementary Metal-Oxide-Semiconductor (CMOS) technology [1, 2, 3, 4, 5] , and superconducting Josephson-junctions [6]. Here, we focus on developing the readout electronics using tunnel-diode oscillator [7, 8] (TDO) circuits. TDO uses a tunnel diode as a negative resistance element [9], generating a microwave signal when connected to an LC circuit. Comparable to cryogenic CMOS devices (typical power consumption: 10 mW [2, 3, 4, 10, 11, 5]) and superconducting Josephson-junction circuits (100 $\mu$[6]), the TDO presented here consumes significantly less power, requiring only 1 $\mu$W.
@@
This research aims to develop a technique that enables scalable qubit read-out schemes for quantum computers operating at cryogenic temperatures. The conventional method, illustrated in Fig. 1(a), is to generate a microwave signal with a room temperature (RT) microwave source and send it to a resonator dispersively coupled to the qubit at the 10 mK stage (also called the mixing chamber, MC stage). The amplitude and phase of the reflected microwave signal depend on the qubit state, allowing us to determine the qubit state. This method is commonly used for qubits made from electrons in semiconductors [12, 13, 14, 15, 16] and a similar method has been proposed for electrons on helium [17].
@@
Figure 1: (a, b) Schematic diagrams illustrating dispersive qubit readout using a resonator: the conventional method in (a) and the tunnel-diode oscillator (TDO) method in (b). The bandwidth of the twisted pair of DC lines is limited to 1 MHz but is sufficient for transmitting the final calculation results to RT and powering the TDO from RT. (c) Schematic of the setup used to characterize the TDO. The characterization includes data measured both with and without the amplifier at 4 K.
@@
In the conventional method, since the input signal is generated at RT, attenuators are placed at the 4 K stage to reduce thermal noise before it reaches the qubit. To prevent the reflected output signal from being overwhelmed by thermal noise at RT, it must be amplified at cryogenic temperatures. Cryogenic amplifiers typically have a lower noise figure, resulting in a higher signal-to-noise ratio. The resonant frequency of the resonator is typically in the range of 100 MHz to 2 GHz [18, 19, 15], and such high-frequency signals are transmitted through coaxial cables. A major scalability issue arises as the number of qubits increases because each qubit requires its own coaxial cable, which are relatively thick (feedthrough connector diameter: $\sim$1 cm, cable diameter: $\sim$1 mm) with attenuators and amplifiers also taking up the limited space inside a cryogenic refrigerator (diameter $\sim$10-100 cm). Multiplexing can reduce the number of coaxial cables to some extent, but is limited by the resonator bandwidth [20].
[…]
@@
In this work, we focus on the characterization of the TDO without utilizing an ADC or a resonator coupled to qubits (Fig. 1(c)). This initial step allows us to evaluate the performance and stability of the TDO as a standalone device under cryogenic conditions, providing a foundation for future integration with qubits. For qubit applications, microwaves are used for qubit state readout and manipulation. Amplitude stability is particularly important for readout [15], while phase noise also plays a crucial role in manipulation [23]. Therefore, in this work, we characterized both amplitude stability and phase noise to evaluate the performance of the TDO in this context.
@@ 2 Tunnel Diode Oscillator
We use the commercial BD-6 Ge-backward tunnel diode (American Microsemiconductor, Inc.) because it has the lowest power consumption among the commercially available options. TDOs with backward tunnel diodes demonstrate frequency stability of 0.001 ppm [8, 24]. Leveraging this precision, they have been employed in the past to investigate various physical properties, such as paramagnetic susceptibility in salts [25], penetration depth in superconductors [26, 27], and the melting point of 3He [28]. These results were obtained with long integration times using a frequency counter, whereas qubit readout demands precise measurements within shorter durations. This emphasizes the need to understand the frequency and amplitude stability achievable over shorter integration times. Furthermore, the oscillation frequencies, typically around 10 MHz [25, 29, 26, 27, 30, 31], are relatively low for qubit readout applications. In the early stages of testing qubit readout, it is also desirable to have the capability to externally tune the frequency. To address this issue, we developed a TDO operating at around 140 MHz, utilizing a microfabricated superconducting spiral inductor. This frequency has reached the lower bound among experiments [32] where the electron transition is measured through dispersive coupling to a resonator, making it applicable for qubit readout. By incorporating a varactor diode as a variable capacitor, we achieved a frequency tunability of 10 MHz without altering the amplitude. Additionally, this oscillator might also be used as a low-power, low-phase-noise clock source for cryo-ADCs.
@@
Figure 2: (a) Conceptual circuit illustration of the TDO. The tunnel diode (TD) can be modeled as a negative resistance $-R$ and a capacitance $C_{\mathrm{TD}}$. The input voltage applied to the TD is converted into and output as a microwave signal with a frequency of $\sim$140 MHz. (b) A photograph of the developed circuit. The pink box encloses the TDO, and the green box the resonator. The orange circle highlights a microfabricated 15-turn spiral inductor on a sapphire substrate, positioned on top of a 1.5-turn pickup coil, enabling a 10% signal to be extracted from the TDO in terms of voltage. The pickup coil and the TDO are connected via a coupler (ZEDC-15). In this study, the resonator is detuned from its resonance, behaving as an open-ended element in the circuit. An enlarged photograph of the microfabricated 15-turn spiral inductor on a sapphire substrate is also shown. The diameter of its boding pad is 176 $\mu$m. (c) Circuit diagram of the device shown in (a). See the main text for details of the circuit components. The signal output from the TDO is measured at RT using a digital oscilloscope or a spectrum analyzer via AC port I or AC port II. For comparison, a signal from a commercial microwave source is sent to the cryogenic device through AC port II, and the reflected signal from the resonator is measured at RT using a digital oscilloscope or a spectrum analyzer via AC port I. The resonator is shaded in gray because its resonant frequency is set far outside the TDO’s operating range, allowing it to function as a high-impedance element and remain inactive. (d) The I-V curve of the tunnel diode BD-6 at the mixing chamber stage temperature (MCT) of 22 mK and 3.4 K. Hysteresis is observed only during oscillation, depending on the sweep direction of the DC bias $V_{\mathrm{TD}}$ applied to the tunnel diode (indicated by arrows). When the bias is increased in the forward direction, oscillation occurs in the range of 0.04 V to 0.245 V, whereas in the reverse direction, the oscillation range is from 0.15 V to 0.04 V.
Fig. 2(a) shows a conceptual circuit illustration explaining how a tunnel diode (TD) implements oscillations. When a TD is connected in an LC resonant circuit, the negative resistance (represented by $-R$ in Fig. 2(a)) cancels out the resistive losses, allowing sustained oscillations. The oscillation frequency is determined by the LC resonance condition:
$f=\frac{1}{2\pi\sqrt{L(C+C_{\mathrm{TD}})}}$ (1)
where $C_{\mathrm{TD}}$ represents the capacitance of the TD. $C$ is the sum of the parasitic capacitance of the circuit board and the capacitance of the varactor diode, and its capacitance value is tunable by changing the voltage applied to the varactor diode, $V_{\mathrm{VD}}$. To achieve high-frequency oscillations, minimizing parasitic capacitance in both the circuit and its components is crucial. Additionally, a high-Q inductor is important, as the self-resonant frequency should be higher than the oscillation frequency.
It is important to carefully select suitable BD-6 tunnel diodes for two reasons: first, to ensure that they exhibit negative resistance at low temperatures, and second, to minimize parasitic capacitance. Although the I-V characteristics measured at RT show little variation among different BD-6 tunnel diodes, some lose their negative resistance at low temperatures (see Appendix A). Therefore, screening using low-temperature I-V measurements is necessary. BD-6 tunnel diodes also exhibit significant variation in parasitic capacitance. We measured 62 different BD-6 tunnel diodes and found that their parasitic capacitance at RT ranged from $C_{\mathrm{TD}}=1.3$ pF to 29.4 pF. The tunnel diode used in this study has a parasitic capacitance of approximately $C_{\mathrm{TD}}=$7 pF (Table. 1).
[…]
@@
Fig. 2(b,c) shows the details of the developed circuit. This circuit board is thermally anchored to the 10 mK stage, which is referred to as the Mixing Chamber (MC) stage. The part enclosed by the pink line represents the TDO, which consists of the Nb spiral inductor and a variable capacitance formed by a varactor diode (MA46H201, MACOM). The varactor diode allows tuning of the LC resonator’s capacitance, thereby varying the oscillation frequency of the TDO. The varactor diode is biased via DC port 2 using a DC voltage source (Yokogawa GS200). The oscillation signal is coupled to the pickup coil with a coupling ratio of 15:1.5. In previous work [25, 29, 26, 27, 30, 31], the signal was extracted capacitively from the same line used to provide the DC bias to the tunnel diode. Instead, in this work, the signal is extracted inductively. This approach ensures that the DC bias line is used exclusively for providing the DC bias, thereby preventing the TDO oscillation from being affected by filters and other components further along the line. As a result, a more stable DC bias can be supplied to the tunnel diode. The tunnel diode is biased via DC port 1 using either the Yokogawa GS200 or a lead-acid battery. After the pickup coil, the signal is sent to AC port II and measured at RT. It is also routed through a directional coupler (ZEDC-15, Mini-Circuits). The coupling from the coupling port to the input/output port of the ZEDC-15 is 20 dB, as measured separately at the 4 K stage. The signal from the coupler is amplified at the 4 K stage using a cryogenic amplifier (CMT-BA1, Cosmic Microwave Technology) and measured at RT with a spectrum analyzer or a digital oscilloscope.
@@
The part enclosed by the green line in Fig. 2(b, c) represents the resonator, which can be coupled to a qubit. In this work, the resonator’s resonant frequency is set far outside the TDO’s operating range, allowing it to act as a high-impedance element and remain unused.
@@
The operating point of the tunnel diode BD-6, where it exhibits negative resistance and oscillates, is approximately $\,\text{V}$ 0.1 V and $I\approx 10\,\mu\text{A}$ (Fig. 2(d)), resulting in a power consumption of about $1\,\mu\text{W}$. When oscillations occur, the DC bias conditions are slightly modified due to the self-rectification of alternating currents by the TD. This explains the hysteresis observed in the I-V curve in Fig. 2(d) depending on the sweep direction of $V_{\mathrm{TD}}$  [24].
@@ 3 Frequency and power tunability
In this section, we demonstrate that the output power and frequency can be controlled by varying the voltage applied to the TD and that the frequency can be independently adjusted by varying the voltage applied to the varactor diode. The frequency variation due to $V_{\mathrm{TD}}$ is shown in Fig. 3(a). The capacitance of the TD, $C_{\mathrm{TD}}$, is voltage-dependent because the distance of the depletion layer in the p-n junction changes as a function of $V_{\mathrm{TD}}$. It is given by [7]
$C_{\mathrm{TD}}=C_{0}\left(1-V_{\mathrm{TD}}/V_{\mathrm{d}}\right)^{-1/2},$ (2)
@@
where $V_{d}=0.5$ V is the diffusion potential (estimated from a measurement) and $C_{0}$ is the capacitance of the TD when $V_{\mathrm{TD}}=0$. Eq. 1 with Eq. 2 has been fitted, as shown by the dashed lines in Fig. 3(a). The capacitance slightly decreases when the MCT is lowered from 3.4 K to 11 mK. The fitting results yield $C=5.8$ pF and $C_{0}=5.7$ pF at MCT = 11 mK, and $C=5.9$ pF and $C_{0}=5.8$ pF at MCT = 3.4 K. Fig. 3(b) demonstrates that the output power can be tuned by approximately 10 dB by adjusting $V_{\mathrm{TD}}$, as the operating point of the TD shifts along its I-V curve. The microwave power that would be delivered to the resonator is around -90 dBm, which is comparable to the optimized power range for qubit readout of electrons used as qubits in semiconductors [34].
Fig. 3(c) shows the frequency tunability as a function of $V_{\mathrm{VD}}$. Considering the additional changes due to $V_{\mathrm{TD}}$, we observe a total frequency tunability of 10 MHz. From this, we estimate that $C$ varies from 7.9 pF to 5.3 pF as $V_{\mathrm{VD}}$ is changed from -1.5 V to 5 V (Table. 1).
@@
Fig. 3(d) shows that for $V_{\mathrm{VD}}>-1.3$ V, the output power remains unaffected by $V_{\mathrm{VD}}$, indicating that the resistive component of the varactor remains unchanged (leakage current stays significantly small within this voltage range as seen in Fig. 7). Fig. 3(d) shows that for $V_{\mathrm{VD}}>-1.3$ V, the output power remains unaffected by $V_{\mathrm{VD}}$, indicating that the resistive component of the varactor remains unchanged. This is consistent with the fact that the leakage current was measured to be significantly low within this voltage range, as shown in Fig. 7. Please note that we have observed a change in output power when $V_{\mathrm{VD}}$ is varied within the same voltage range at RT when using a similar circuit with a normal coil inductor. We attribute this difference to the fact that the leakage current becomes significantly suppressed at low temperatures.
@@
Figure 3: The signal from the TDO obtained through AC port II was measured using a spectrum analyzer. (a,b) $V_{\mathrm{VD}}$ is set to 0. The sweep direction of $V_{\mathrm{TD}}$ is forward. (a) Output frequency as a function of $V_{\mathrm{TD}}$ at MCT = 11 mK and 3.4 K. The dashed lines represent the fitted curves, which account for changes in depletion layer thickness induced by varying $V_{\mathrm{TD}}$. See the main text for details on the fitting procedure. (b) Output power as a function of $V_{\mathrm{TD}}$ at MCT = 11 mK and 3.4 K. The difference in background levels is attributed to the use of different spectrum analyzers for measurements at MCT = 11 mK and 3.4 K. (c) Output frequency as a function of $V_{\mathrm{VD}}$ at MCT = 11 mK. (d) Output power as a function of $V_{\mathrm{VD}}$ at MCT = 11 mK. It remains constant for $V_{\mathrm{VD}}>-1.3$ V.
Table 1: The table presents capacitance values under different voltage conditions at MCT = 11 mK. At 3.4 K, the capacitance values are higher by approximately 0.1 pF compared to those at 11 mK in all cases.
$C_{\text{TD}}$ $C$
Voltage (V) $V_{\text{TD}}=0$ $V_{\text{TD}}=0.08$ $V_{\text{TD}}=0.18$ $V_{\text{VD}}=-1.5$ $V_{\text{VD}}=0$ $V_{\text{VD}}=5$
Capacitance (pF) 5.7 6.3 7.2 7.9 5.8 5.3